OpenStudy (anonymous):

Use the distributive property to multiply. 12(u + v)

Parth (parthkohli):

Distributive property says \(a(b + c) = ab + ac\)

Parth (parthkohli):

Replace a with 12, b with u and c with v.

Parth (parthkohli):

\( \color{Black}{\Rightarrow 12(u + v) = 12u + 12v }\)